An arrow is shot vertically upward from a platform 40 feet high at a rate of 224 ft/s. This is modeled by the equation h=-16t^2+224t+40. When will the arrow hit the ground? (Round to three decimal places) Answer by ankor@dixie-net.com(22740) 
 An arrow is shot vertically upward from a platform 40 feet high at a rate of 224 ft/s.
This is modeled by the equation h=-16t^2+224t+40.
When will the arrow hit the ground? (Round to three decimal places)
:
When the arrow hits the ground, h = 0, therefore
-16t^2 + 224t + 40 = 0
use the quadratic formula to find t; a=-16; b=224; c=40
The reasonable solution
t = 14.176 seconds
:
:
Check this using you calc with t=14.176
-16(14.176^2) + 224(14.176 + 40 =